### v1

@@ Abstract
Do vision-language models (VLMs) pre-trained to caption an image of a durian learn visual concepts such as brown (color) and spiky (texture) at the same time? We aim to answer this question as visual concepts learned “for free” would enable wide applications such as neuro-symbolic reasoning or human-interpretable object classification. We assume that the visual concepts, if captured by pre-trained VLMs, can be extracted by their vision-language interface with text-based concept prompts. We observe that recent works prompting VLMs with concepts often differ in their strategies to define and evaluate the visual concepts, leading to conflicting conclusions. We propose a new concept definition strategy based on two observations: First, certain concept prompts include shortcuts that recognize correct concepts for wrong reasons; Second, multimodal information (e.g. visual discriminativeness, and textual knowledge) should be leveraged when selecting the concepts. Our proposed concept discovery and learning (CDL) framework is thus designed to identify a diverse list of generic visual concepts (e.g. spiky as opposed to spiky durian), which are ranked and selected based on visual and language mutual information. We carefully design quantitative and human evaluations of the discovered concepts on six diverse visual recognition datasets, which confirm that pre-trained VLMs do learn visual concepts that provide accurate and thorough descriptions for the recognized objects. All code and models are publicly released. [^1] Code is available at: https://github.com/brown-palm/Concept-Discovery-and-Learning [^2] Project homepage: https://conceptdiscovery.github.io
Keywords:
Visual Concepts Vision-Language Models
@@ 1 Introduction
Can a vision-language model (VLM) pre-trained on images of California seagull learn visual concepts [13, 12, 14, 28], such as yellow legs and white belly, to describe an image of a Gentoo penguin, which the model might not see during training? Visual concepts such as color, shape, and texture help models generalize compositionally [7, 22, 9, 29], and can be incorporated into neuro-symbolic frameworks [19, 37] or offer concept-based explanations for classification decisions [10].
@@
Our paper aims to investigate if VLMs, such as CLIP [25], learn visual concepts automatically when pre-trained on image and text pairs with contrastive learning objectives. We hypothesize that the visual concepts, if captured by the pre-trained VLMs, can be directly extracted by prompting their vision-language interface, without needing to probe their internal representations. Our research question can thus be formulated as discovering the visual concept prompts, and evaluating the quality of the extracted concepts. Intuitively, the desirable visual concepts should be precise, to faithfully reflect the visual concepts presented in an image; They should also be thorough, so that the concepts can provide most of the distinctive visual features for an object of interest.
@@
Figure 1: The design of concept prompts plays a critical role on the understanding of whether VLMs learn visual concepts. We can observe that concept-augumented prompt can predict correct correct visual concepts (e.g., yellow ring around black eyes) when the prompt is associated with the category name (California seagull). When the category name is removed from the prompt (second column), the retrieved concepts are either non-visual or incorrect. We attribute this to the category name bias (third column), as the correct category can be retrieved by CLIP even when the paired descriptions are randomly shuffled and thus irrelevant. We propose a concept discovery and learning (CDL) framework and demonstrates that pre-trained VLMs can indeed learn visual concepts (e.g., last column). Correct predictions and concepts are in green, while wrong concepts are in red and non-visual concepts are in violet. The category names are in orange.
Interestingly, several recent works [21, 35, 39] studying visual concepts in pre-trained VLMs reached different conclusions. We attribute these discrepancies to their different strategies for concept prompt selection and evaluation. For example, Yun et al. [39] observed that CLIP does not appear to recognize fine-grained visual attributes for birds [30], when a pre-defined list of visual concepts by bird experts was used. In contrast, Menon et al.  [21] demonstrated that object prompts with concepts proposed by a large language model (LLM) appear to provide interpretable object classification, as the concept descriptions are nicely correlated with the recognized object categories. As illustrated in Figure 1, we observe that the design of concept prompts has a substantial impact on the “recognized” concepts: when certain shortcuts are included (e.g., the object category California seagull) in the prompt, the retrieved concepts (e.g., black eyes) can be correct for wrong reasons, thus undermining their ability to answer our research question. We also observe that the visual concepts need to be selected based on both visual and language cues, as certain concepts returned by an LLM is correct, but cannot be recognized visually (e.g., fish-eating). As for evaluation, proxy benchmarks such as classification accuracy [10] are often used, which may not always positively correlate with the quality of the visual concepts [20].
In response to these observations, we propose a new concept discovery strategy to eliminate the shortcuts and utilize both visual and language information, in order to properly investigate the concepts in pre-trained VLMs. We first use a large and diverse image captioning dataset [26] as a source of objects to discover diverse and thorough visual concepts shared by multiple objects. We then rank and select the visual concepts based on multimodal information: Given a collection of images and their text descriptions, we prefer the visual concepts that can be both reliably recognized from the images (e.g., with a pre-trained VLM), and deemed as suitable descriptions based on the text descriptions (e.g., according to the prior knowledge encoded by an LLM). We compute this ranking with a mutual information formulation. When evaluating the concepts on a specific domain, we can use the same formulation to select a compact list of concepts. Finally, we propose a self-supervised method to adjust the final linear projection layer of pre-trained VLMs, which we show empirically can further improve the quality of the concepts without fine-tuning the encoder backbones. We name our overall framework as Concept Discovery and Learning (CDL).
We observe that as expected, when applying the concepts extracted by CDL for interpretable object recognition, our method consistently outperforms baselines in both full-shot and few-shot settings, across six visually diverse benchmarks. More importantly, we introduce a suite of quantitative and human evaluation protocols to measure the precision and thoroughness of concepts extracted by CDL. Across different evaluation protocols, we observe that the concepts extracted by CDL from CLIP are both precise and thorough, demonstrating that pre-trained vision-language models can indeed learn visual concepts via its vision-language interface. Our models and code will be publicly released.
@@ 2 Related Work
Vision-and-language models (VLMs) pretrained on unlabeled pairs of images and texts from the internet have shown great success on multi-modal benchmarks. Representations learned by these VLMs can be transferred to a wide range of tasks, such as visual question answering [17, 16, 2], and image and video captioning [18, 40, 34]. These pretrained VLMs can directly recognize complex concepts such as object categories in a zero-shot setting with text prompts [1, 25]. However, it remains unclear whether VLMs recognize composite concepts by learning and reasoning over primitive concepts such as colors and shapes. Understanding the roles of these visual concepts is important since they can serve as clues for interpretable fine-grained visual recognition [10, 8, 4], zero-shot or few-shot learning [21, 24], open-world object detection [36] and visual reasoning [19].
@@
Previous studies have explored the capability of VLMs to capture and compose primitive concepts [39, 38] and bind visual concepts with objects [15, 31]. Yun et al. [39] demonstrate that VLMs do not capture composable primitive concepts by intervening a linear classifier learned from VLM-predicted concepts. Meanwhile, previous works [24, 21] show that concept-augmented prompts do provide improvements for VLM-based image recognition. In this paper, we aim to address this discrepency and understand pretrained VLMs’ true capability of encoding interpretable visual concepts.
Previous research [10] have proposed the Concept Bottleneck Models (CBM) to decompose the end-to-end decision into concept-level reasoning. The concept bottlenecks can provide interpretable decision basis for complex recognition and reasoning, and are applied in various machine learning domains such as medical diagnosis [33]. Prior works [35, 32] have proposed to build the CBM with descriptive concepts to improve the recognition performance of VLMs.
@@ 3 Method
We first describe a framework of extracting visual concepts from pre-trained VLMs and its application for object recognition in Section 3.1. We show in Section 3.2 that the text prompts used to extract visual concepts may include shortcuts, and the concept activations may not correspond to how likely the visual concepts actually present in an image. We address this issue from two perspectives: We first propose a new framework in Sections 3.3 and 3.4 to discover the visual concepts that are category-generic and visually salient from a pre-trained VLM. We then propose a suite of quantitative and human evaluations on the interpretability, precision, thoroughness, and generalizability of the discovered visual concepts in Section 3.5. Together they help us understand if pre-trained VLMs learn to encode visual concepts.
@@ 3.1 Object Recognition via Visual Concepts
Vision-language models such as CLIP [25] jointly learn an image encoder $\mathcal{E_{I}}$ and a text encoder $\mathcal{E_{T}}$ to align images and texts in a shared embedding space. Thanks to the flexibility of the image-language interface, several recent works [24, 21, 39, 35, 32] attempted to construct semantically interpretable representations by projecting an encoded visual embedding with basis defined by encoded text embeddings. The text embeddings are obtained by encoding manually designed or automatically generated text “prompts” that are likely to correspond to visual concepts. Concretely, given an image $I$ and a set of concept prompts $\mathbb{P}$ of $N$ concepts, one can project the visual features $\mathcal{E_{I}}(I)$ into the space of concepts as an $n$-dimensional concept activation vector $\mathbf{a}=[a_{1},a_{2},...,a_{N}]$. Each concept activation is computed as $a_{i}=\text{sim}(\mathcal{E_{I}}(I),\mathcal{E_{T}}(p_{i}))$, where $p_{i}$ is the $i$-th concept prompt, and $\text{sim}(\cdot)$ is a similarity function, such as cosine similarity, that measures the similarity between the encoded visual and text embeddings. Our paper assumes that the visual concepts, if well learned by a VLM, can be extracted as concept activations $\mathbf{a}$ via the text prompts.
The concept activations can be utilized for multi-modal object recognition with a function $f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{M}$ that predicts object categories, where $M$ is the total number of categories. When $f$ is a linear function $f(\mathbf{a})=\mathbf{a}\cdot\mathcal{W}$, the learned $N\times M$ weight matrix $\mathcal{W}$ allows us to interpret how the visual concepts are utilized for object recognition. A higher positive weight $w_{ij}$ indicates the $i$-th concept is deemed as important positive evidence for recognizing the $j$-th object category, and a near-zero weight indicates the concept is deemed as irrelevant. The linear classifier that maps concept activations to categorical predictions is referred to as Concept Bottleneck Models [10] (CBM), which have been adopted to study concept learning with VLMs in [35, 39, 32]. For zero-shot object recognition when $f(\cdot)$ cannot be learned from data, some works [24, 21] assume the visual concepts are category-specific (hence the object names are included in the prompts), and simplify $f(\cdot)$ to be a linear function $f_{j}(\mathbf{a})=\frac{1}{|\mathcal{C}_{j}|}\sum_{a_{i}\in\mathcal{C}_{j}}a_{i}$, where $\mathcal{C}_{j}$ is the set of visual concepts for the $j$-th category, and $\mathcal{C}_{j}\cap\mathcal{C}_{k}=\emptyset$. We observe that this assumption does not hold for real-world datasets as many visual concepts are often shared by several objects.
@@ 3.2 Do Prompted Activations Correspond to Visual Concepts?
Large language models (LLMs) are often utilized as the knowledge source to propose the relevant visual concept prompts given an object category [24, 21, 35, 32]. One example is illustrated in Figure 1, where for California Gull, concept prompts such as “California Gull, which has a long, black tail” are proposed. As discussed in Section 3.1, the linear function $f(\mathbf{a})$ allows us to identify the most important visual concepts for object recognition by picking the highest weighted $w_{ij}$ concepts $i$’s for object $j$. One can then consider two proxy evaluations to measure whether the prompted concept activation $a_{i}$ actually corresponds to the visual concept $i$ that is present in an image: First, by measuring the object classification accuracy, and assuming the higher the accuracy is, the more precise the concept activations are. Second, by comparing the top ranked concepts for an object category and those identified by human experts, which allows us to understand not only the precision but also the thoroughness of the concepts.
@@
We observe that these proxy evaluations, while intuitive, require careful design of concept prompting strategy in order to draw conclusions on whether the concepts are actually encoded by the pre-trained VLMs. The first issue we identify is the existence of certain shorts in the text prompts, leading to “false positive” conclusions. For example, Figure 1 illustrates the concepts with the highest activations according to CLIP. Although the first column appears to indicate that most of the selected concepts are semantically correlated with the input image of a California Gull, it remains unclear whether the concepts are retrieved because they are recognized by CLIP or the class name is utilized as a shortcut. We perform a simple ablation to investigate this potential shortcut: In the second column of Figure 1, we observe that irrelevant concepts are selected when category names are removed from the prompts. In the third column, we observe that when we combine the category names with randomly shuffled descriptors, CLIP tends to align images with concepts that contain correct category names and wrong descriptors. This phenomenon indicates that the category names, instead of the visual concepts, are used to generate the concept activations. We reinforce these qualitative observations in Section 4.2, where we demonstrate consistently across six datasets that the zero-shot classification accuracy remains similar when LLM-generated concepts or randomly shuffled concepts are paired with category names, and that the accuracy drops significantly when category names are removed from the text prompts.
@@
A second issue is the existence of sub-optimal text prompts, leading to “false-negative” conclusions. For example, the concepts marked with purple background in Figure 1 are not visually recognizable (e.g. “nocturnal bird”), and hence should not be considered as visual concepts. Similarly, we observe that the text prompts used by Yun et al. [39] are designed by birding experts, which may not be in a friendly format to serve as text prompts (e.g. “crested head pattern”).

### v2

@@ Abstract
Do vision-language models (VLMs) pre-trained to caption an image of a durian learn visual concepts such as brown (color) and spiky (texture) at the same time? We aim to answer this question as visual concepts learned “for free” would enable wide applications such as neuro-symbolic reasoning or human-interpretable object classification. We assume that the visual concepts, if captured by pre-trained VLMs, can be extracted by their vision-language interface with text-based concept prompts. We observe that recent works prompting VLMs with concepts often differ in their strategies to define and evaluate the visual concepts, leading to conflicting conclusions. We propose a new concept definition strategy based on two observations: First, certain concept prompts include shortcuts that recognize correct concepts for wrong reasons; Second, multimodal information (e.g. visual discriminativeness, and textual knowledge) should be leveraged when selecting the concepts. Our proposed concept discovery and learning (CDL) framework is thus designed to identify a diverse list of generic visual concepts (e.g. spiky as opposed to spiky durian), which are ranked and selected based on visual and language mutual information. We carefully design quantitative and human evaluations of the discovered concepts on nine diverse visual recognition datasets, which confirm that pre-trained VLMs do learn visual concepts that provide accurate and thorough descriptions for the recognized objects. Code and models are publicly released.  [^1] Project page and code: https://conceptdiscovery.github.io
@@ 1 Introduction
Can a vision-language model (VLM) pre-trained on images of California seagull learn visual concepts (Lake et al., 2015; 2011; Lee et al., 2023; Sun et al., 2015), such as yellow legs and white belly, to describe an image of a Gentoo penguin, which the model might not see during training? Visual concepts such as color, shape, and texture help models generalize compositionally (Farhadi et al., 2009; Nagarajan & Grauman, 2018; Hsieh et al., 2024; Thrush et al., 2022; Stein et al., 2024), and can be incorporated into neuro-symbolic frameworks (Mao et al., 2019; Yi et al., 2018) or offer concept-based explanations for classification decisions (Koh et al., 2020).
@@
Our paper aims to investigate if VLMs, such as CLIP (Radford et al., 2021), learn visual concepts automatically when pre-trained on image and text pairs with contrastive learning objectives. We hypothesize that the visual concepts, if captured by the pre-trained VLMs, can be directly extracted by prompting their vision-language interface, without needing to probe their internal representations. Our research question can thus be formulated as discovering the visual concepts encoded by pre-trained VLMs, and evaluating the quality of the extracted concepts.
@@
Figure 1: The top-selected text prompts by CLIP given a query image, VDES is proposed by Menon & Vondrick (2022), whereas CDL is our proposed approach. The design of concept prompts plays a critical role on understanding whether VLMs learn visual concepts. We can observe that concept-augumented prompt can predict correct visual concepts (e.g., gray back and wings) when the prompt is associated with the category name (California seagull). When the category name is removed from the prompt (Column 2), the retrieved concepts are either non-visual or incorrect. We attribute this to the category name bias (Column 3), as the correct category can be retrieved by CLIP even when the paired descriptions are randomly shuffled and thus irrelevant. We propose a concept discovery and learning (CDL) framework and demonstrate that pre-trained VLMs can indeed learn visual concepts (e.g., Column 4). Correctly predicted concepts are in green, wrong concepts are in red, and non-visual concepts are in violet. Category names are in orange.
Interestingly, several recent works (Menon & Vondrick, 2022; Yang et al., 2023b; Yun et al., 2023) reached different conclusions on the encoding of visual concepts in pre-trained VLMs. For example, Yun et al. (2023) observed that CLIP does not appear to recognize fine-grained visual attributes for birds (Wah et al., 2011), where the list of visual attributes is pre-defined by bird experts. In contrast,  Menon & Vondrick (2022) demonstrated that object prompts with visual concepts proposed by a large language model (LLM) appear to provide interpretable object classification, as the concept descriptions are nicely correlated with the recognized object categories. As illustrated in Figure 1, we attribute these discrepancies to their different strategies for extracting visual concepts in a pre-trained VLM: First, according to the first and second columns of Figure 1, certain shortcuts (e.g., the object category California seagull) may dominate the text prompts when recognizing visual concepts, reinforcing the object-concept correlations as given by a prior knowledge base, such as LLMs. A possible remedy is to discover general visual concepts shared by multiple objects, without including the shortcuts in a text prompt. Second, according to the third column in Figure 1, some of concepts proposed by LLMs are not visually discriminative (e.g., “a nocturnal bird”) or cannot be reliably recognized by VLMs (e.g. concept prompts marked in red, or the expert prompts used by Yun et al. (2023)). They should be excluded from the list of visual concepts. More generally, in order to draw a convincing conclusion on whether pre-trained VLMs learn to encode visual concepts, one should discover the list of visually discriminative concepts and use them to prompt VLMs in a shortcut-free manner, and then measure the quality and thoroughness of the discovered concepts, when they are tasked to recognize fine-grained objects and generate interpretable explanations with visual concepts.
In this paper, we propose a novel concept discovery and learning framework that extracts shortcut-free and visually discriminative concepts from pre-trained VLMs. To discover general concepts that do not include category specific shortcuts, we propose to utilize a large and diverse image captioning dataset to discover general concepts shared by multiple objects. To select concepts that are well-recognizable by VLMs, we propose to jointly consider LLM and VLM information during concept discovery. We prefer the visual concepts that can be both reliably recognized from the images by a VLM, and deemed as suitable descriptions for the objects of interest in the images by an LLM. Furthermore, although visual concepts can be discovered from frozen pre-trained VLMs, they are not always perfectly aligned with images (see Column 2 in Figure 1). To improve the quality of selected concepts, we propose a self-supervised method to fine-tune the last projection layer of VLMs by leveraging the knowledge already encoded in pre-trained VLMs and LLMs. We name our overall framework as Concept Discovery and Learning (CDL).
We design a suite of evaluation protocols for the quality of the extracted concepts. Intuitively, the desirable visual concepts should be precise, to faithfully reflect the visual concepts presented in an image; They should also be thorough, so that the concepts can provide most of the distinctive visual features for an object of interest. Furthermore, these concepts should be generalizable, to effectively describe new and unseen objects. According to the experimental results, the proposed CDL framework effectively discover and learn precise, thorough and generalizable concepts. The discovered concepts lead to accurate and interpretable object recognition on nine visually diverse benchmarks. The results demonstrate that pre-trained VLMs can indeed learn visual concepts via its vision-language interface, and these visual concepts can lead to performant and interpretable visual recognition.
@@ 2 Related Work
Vision-and-language models (VLMs) pretrained on unlabeled pairs of images and texts from the internet have shown great success on multi-modal benchmarks. Based on the pre-training objectives, VLMs can be broadly categorized into two primary types: contrastive VLMs and generative VLMs. Contrastive VLMs (Radford et al., 2021; Alayrac et al., 2022; Jia et al., 2021; Yao et al., 2021) focus on learning joint representations of images and texts by leveraging contrastive learning techniques (Chen et al., 2020). Generative VLMs Lu et al. (2019); Wang et al. (2023); Chen et al. (2023); Liu et al. (2024) aim to generate coherent text descriptions from images. Representations learned by these VLMs can be transferred to a wide range of tasks, such as image classification (Pratt et al., 2023), visual question answering (Li et al., 2023; Bai et al., 2023), and image and video captioning (Zhang et al., 2021; Yang et al., 2023a).
@@
Visual concepts, which represent the fundamental factors of variations (e.g. colors, shapes) (Hu et al., 2018) in the visual world, have been widely utilized to develop intepretable visual models that can generalize compositionally (Farhadi et al., 2009; Lampert et al., 2009; Nagarajan & Grauman, 2018; Stein et al., 2024). Previous works (Sun et al., 2015; Hernandez et al., 2021; Lee et al., 2023) have proposed to discover textual terms associated to visual concepts to interpret visual systems. These concepts can be integrated into neuro-symbolic frameworks (Mao et al., 2019; Yi et al., 2018) and offer concept-based explanations for visual recognition (Koh et al., 2020). Contrastive VLMs, learning joint representation for images and texts, provide a natural interface for discovering such visual concepts. However, it remains unclear whether VLMs pre-trained with contrastive objectives can learn discoverable visual concepts. Previous studies have explored the capability of VLMs to capture and compose primitive concepts (Yun et al., 2023; Yuksekgonul et al., 2022a) and bind visual concepts with objects (Lewis et al., 2022; Yamada et al., 2022). Yun et al. (2023) demonstrate that VLMs do not capture composable primitive concepts by intervening a linear classifier learned from VLM-predicted concepts. Meanwhile, previous works (Pratt et al., 2023; Menon & Vondrick, 2022) show that concept-augmented prompts do provide improvements for VLM-based image recognition. In this paper, we aim to address this discrepency and understand pretrained VLMs’ true capability of encoding interpretable visual concepts.
To interpret the decision of visual models with visual concepts, Koh et al. (2020) proposed Concept Bottleneck Model (CBM) to decompose the visual recognition into concept classification and concept-category mapping. While an alternative approach (Bau et al., 2017; Kim et al., 2018; Hernandez et al., 2021) aims to develop post-hoc interpretation for visual models by directly analyzing the internal representation of the models within the concept space, the CBM does not rely on the model having already learned concepts, instead, it simultaneously trains the concept classifier and linear concept-category mapper upon the model being analyzed with the image classification objective. As a result, the CBM can evaluate the concept learning ability of a model through the concept-category map, regardless of whether the model was previously trained to learn those concepts. Recent works have developed various methods to enhance the performance and adaptability of CBMs (Yuksekgonul et al., 2022b; Hu et al., 2024) and have applied them to machine learning tasks that require high transparency, such as medical diagnosis (Yan et al., 2023b; Chauhan et al., 2023). However, previous CBM-based approaches typically rely on concept classifiers trained with supervision. In this paper, we propose to build CBMs based on the pre-trained VLMs to evaluate their concept learning ability. Assuming that pre-trained VLMs inherently learn visual concepts, we directly extract concept-image similarities from the VLM encoders to serve as the concept classifier and only train a one-layer concept-to-category mapper.
@@ 3 Method
We first describe a framework for extracting visual concepts from pre-trained VLMs and its application for object recognition in Section 3.1. We show in Section 3.2 that concepts which are non-visual or include certain shortcuts might lead to wrong correlations, where the concept activations do not correspond to how likely the visual concepts are actually present in an image. In Section 3.3, we propose a concept discovery method to select shortcut-free visual concepts from a large image captioning dataset by utilizing VLM and LLM knowledge to evaluate the visual discrimination. We also propose a self-supervised learning framework to re-align the image-text interfaces of VLMs to improve the quality of selected concepts. In Section 3.4, we describe the method to select a compact set of concepts for specific object recognition tasks. In Section 3.5 we propose a suite of quantitative and human evaluations on the interpretability, precision, thoroughness, and generalizability of the discovered visual concepts. Together they help us understand if pre-trained VLMs learn to encode visual concepts.
@@ 3.1 Object Recognition via Visual Concepts
Vision-language models such as CLIP (Radford et al., 2021) jointly learn an image encoder $\mathcal{E_{I}}$ and a text encoder $\mathcal{E_{T}}$ to align images and texts in a shared embedding space. Thanks to the flexibility of the image-language interface, several recent works (Pratt et al., 2023; Menon & Vondrick, 2022; Yun et al., 2023; Yang et al., 2023b; Yan et al., 2023a) attempted to construct semantically interpretable representations by projecting an encoded visual embedding with basis defined by encoded text embeddings. The text embeddings are obtained by encoding manually designed or automatically generated text “prompts” that are likely to correspond to visual concepts. Concretely, given an image $I$ and a set of concept prompts $\mathbb{P}$ of $N$ concepts, one can project the visual features $\mathcal{E_{I}}(I)$ into the space of concepts as an $n$-dimensional concept activation vector $\mathbf{a}=[a_{1},a_{2},...,a_{N}]$. Each concept activation is computed as $a_{i}=\text{sim}(\mathcal{E_{I}}(I),\mathcal{E_{T}}(p_{i}))$ (1), where $p_{i}$ is the $i$-th concept prompt, and $\text{sim}(\cdot)$ is a similarity function, such as cosine similarity, that measures the similarity between the encoded visual and text embeddings. The activations are standardized with z-score normalization to ensure comparability across objects. Our paper assumes that the visual concepts, if well learned by a VLM, can be extracted as concept activations $\mathbf{a}$ via the text prompts.
The concept activations can be utilized for multi-modal object recognition with a function $f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{M}$ that predicts object categories, where $M$ is the total number of categories. When $f$ is a linear function $f(\mathbf{a})=\mathbf{a}\cdot\mathcal{W}$, the learned $N\times M$ weight matrix $\mathcal{W}$ allows us to interpret how the visual concepts are utilized for object recognition. A higher positive weight $w_{ij}$ indicates the $i$-th concept is deemed as important positive evidence for recognizing the $j$-th object category, and a near-zero weight indicates the concept is deemed as irrelevant. The linear classifier that maps concept activations to categorical predictions is referred to as Concept Bottleneck Models (Koh et al., 2020) (CBM), which have been adopted to study concept learning with VLMs in Yang et al. (2023b); Yun et al. (2023); Yan et al. (2023a). For zero-shot object recognition when $f(\cdot)$ cannot be learned from data, some works (Pratt et al., 2023; Menon & Vondrick, 2022) assume the visual concepts are category-specific (hence the object names are included in the prompts), and simplify $f(\cdot)$ to be a linear function $f_{j}(\mathbf{a})=\frac{1}{|\mathcal{C}_{j}|}\sum_{a_{i}\in\mathcal{C}_{j}}a_{i}$, where $\mathcal{C}_{j}$ is the set of visual concepts for the $j$-th category, and $\mathcal{C}_{j}\cap\mathcal{C}_{k}=\emptyset$. We observe that this assumption does not hold for real-world datasets as many visual concepts are often shared by several objects, as illustrated in Figure 1.
@@ 3.2 Do Prompted Activations Correspond to Visual Concepts?
Large language models (LLMs) are often utilized as the knowledge source to propose the relevant visual concept prompts given an object category (Pratt et al., 2023; Menon & Vondrick, 2022; Yang et al., 2023b; Yan et al., 2023a). One example is illustrated in Figure 1, where for California Gull, concept prompts such as “California Gull, which has a long, black tail” are proposed. As discussed in Section 3.1, the linear function $f(\mathbf{a})$ allows us to identify the most important visual concepts for object recognition by picking the highest weighted $w_{ij}$ concepts $i$’s for object $j$. One can then consider two proxy evaluations to measure whether the prompted concept activation $a_{i}$ actually corresponds to the visual concept $i$ that is present in an image: First, by measuring the object classification accuracy, and assuming that the higher the accuracy is, the more precise the concept activations are. Second, by comparing the top ranked concepts for an image of a certain object category and those identified by human experts, which allows us to understand not only the precision but also the thoroughness of the concepts.
@@
We observe that these proxy evaluations, while intuitive, require careful design of concept prompting strategy in order to draw conclusions on whether the concepts are actually encoded by the pre-trained VLMs. The first issue we identify is the existence of certain shortcuts in the text prompts, leading to “false positive” conclusions. For example, Figure 1 illustrates the concepts with the highest activations according to CLIP. Although the first column appears to indicate that most of the selected concepts are semantically correlated with the input image of a California Gull, it remains unclear whether the concepts are retrieved because they are recognized by CLIP or the class name is utilized as a shortcut. We perform a simple ablation to investigate this potential shortcut: In the second column, we observe that when we combine the category names with randomly shuffled descriptors, CLIP tends to align images with concepts that contain correct category names and wrong descriptors. This phenomenon indicates that the category names, instead of the visual concepts, are used to generate the concept activations. We validate these qualitative observations in Section 4.2, where we demonstrate consistently across nine datasets that the zero-shot classification accuracy remains similar when LLM-generated concepts or randomly shuffled concepts are paired with category names, and that the accuracy drops significantly when category names are removed from the text prompts.
@@
A second issue is the existence of sub-optimal text prompts, leading to “false-negative” conclusions. For example, the concepts marked with purple background in Figure 1 are not visually recognizable (e.g. “nocturnal bird”), and hence should not be considered as visual concepts. Similarly, we observe that the text prompts used by  Yun et al. (2023) are designed by birding experts, which may not be in a friendly format to serve as text prompts (e.g. “crested head pattern”).